# Fast Penalized Generalized Estimating Equations for Large Longitudinal Functional Datasets

### v1

@@ 2.2 One-step fGEE
To scale fGEE to large datasets, we propose a one-step estimator of the form
$\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)}=\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}-\left[\widehat{\mathbb{E}}\left(\nabla_{\bm{\theta}}\bm{U}_{\Lambda_{1}}(\mathbf{X}_{i},\bm{Y}_{i};\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})\right)\right]^{-1}\frac{1}{N}\sum_{i=1}^{N}\bm{U}_{\Lambda_{1}}\left(\mathbf{X}_{i},\bm{Y}_{i};\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}\right),$ (3)
where $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$ is an initial estimate fit with smoothing parameters $\Lambda_{0}$. We estimate (3) as
@@
$\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)}=\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}+\left[\frac{1}{N}\sum_{i=1}^{N}[\mathbb{D}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})]^{T}[\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})]^{-1}[\mathbb{D}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})]+\Lambda_{1}\mathbb{S}\right]^{-1}\times~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
$\frac{1}{N}\sum_{i=1}^{N}\left\{[\mathbb{D}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})]^{T}[\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})]^{-1}\left[\bm{Y}_{i}-\widehat{\bm{\mu}}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})\right]-\Lambda_{1}\mathbb{S}\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}\right\},$ (4)
@@
where we plug in $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$ to calculate $\mathbb{D}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})$, $\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})$, and $\widehat{\bm{\mu}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})=g^{-1}(\mathbb{X}_{i}\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})$, with $g^{-1}$ applied component-wise. The updated estimate for the functional coefficient of covariate $r$ is then obtained as $\widehat{\bm{\beta}}_{\Lambda_{1},r}^{(1)}=\mathbf{B}\widehat{\bm{\theta}}_{\Lambda_{1},r}^{(1)}$. Any consistent estimator for $\bm{\theta}$ can be used for $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$; we use (2) with a working independence correlation structure, i.e., a function-on-scalar regression that ignores within-cluster correlation of outcome values across $s$ and $j$, and uses the same penalty and spline bases as the fGEE. In practice, we estimate $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$ with the pffr function (Scheipl et al.,, 2015) from the refund package (see Appendix A.3 for details). We formalize the necessary consistency properties of $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$ for the population parameter, $\bm{\theta}$ in Section 3.
@@
The one-step can be conceptualized as a de-biasing of, or equivalently a single Newton-Raphson step from, the initial estimate $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$. It is much faster than the fully-iterated fGEE, because it requires inversion of a working covariance matrix only twice per cluster: 1) $[\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})]^{-1}$ to estimate $\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)}$, and 2) $[\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)})]^{-1}$ to estimate $\widehat{\text{Var}}\left(\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)}\right)$ (see Section 2.5).
2.3 Working Correlations

### v2

@@ 2.2 One-step fGEE
To scale fGEE to large datasets, we propose a one-step estimator of the form $\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)}=\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}-\left\{\widehat{\mathbb{E}}\left(\nabla_{\bm{\theta}}\bm{U}_{\Lambda_{1}}(\mathbf{X}_{i},\bm{Y}_{i};\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})\right)\right\}^{-1}\frac{1}{N}\sum_{i=1}^{N}\bm{U}_{\Lambda_{1}}\left(\mathbf{X}_{i},\bm{Y}_{i};\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}\right)$, where $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$ is an initial estimate fit with smoothing parameters $\Lambda_{0}$. We denote $\mathbb{D}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})=\frac{\partial\bm{\mu_{i}}(\bm{\theta})}{\partial\bm{\theta}}\big|_{\bm{\theta}=\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}}}$, $\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})$ as the working covariance estimated with $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$, and $\widehat{\bm{\mu}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})=g^{-1}(\mathbb{X}_{i}\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})$ with $g^{-1}$ applied component-wise. Our proposed estimator is formulated as follows
@@
$\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)}=\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}+\left[\frac{1}{N}\sum_{i=1}^{N}\{\mathbb{D}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})\}^{T}\{\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})\}^{-1}\{\mathbb{D}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})\}+\Lambda_{1}\mathbb{S}\right]^{-1}\times$~~~~~~~~~~
$\frac{1}{N}\sum_{i=1}^{N}\left[\{\mathbb{D}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})\}^{T}\{\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})\}^{-1}\left\{\bm{Y}_{i}-\widehat{\bm{\mu}}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})\right\}-\Lambda_{1}\mathbb{S}\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}\right].$ (3)
@@
Any consistent estimator for $\bm{\theta}$ can be used for $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$; we use a parametric function-on-scalar regression that ignores within-cluster correlation of outcome values across $s$ and $j$, and uses the same penalty and spline bases as the fGEE (similar to solving (2) with a working independence correlation). In practice, we estimate $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$ with the refund::pffr() function (Scheipl et al., 2016) (see Appendix A.9 for details). We formalize the necessary consistency properties of $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$ for the population parameter, $\bm{\theta}$ in Section 3.
@@
The one-step can be conceptualized as a de-biasing of, or equivalently a single Newton-Raphson step from, the initial estimate $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$. It is much faster than the fully-iterated fGEE, because it requires inversion of a working covariance matrix only twice per cluster: 1) $\{\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})\}^{-1}$ to estimate $\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)}$, and 2) $\{\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)})\}^{-1}$ to estimate $\widehat{\text{Var}}\left(\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)}\right)$ (see Section 2.2). We summarize the estimation steps as an algorithm in Appendix S1.
Working Correlations